Given that X and Y are independent random variables and E(X) = 4, E(Y) = 7, Var(X) = 3, Var(Y) = 6; find Var(5X - 6).


1
Expert's answer
2022-01-18T09:14:31-0500

Solution:

Given, E(X) = 4, E(Y) = 7, Var(X) = 3, Var(Y) = 6

"Var(5X - 6)=25Var(X)" "[\because Var[aX + b] = a^2Var(X)]"

"\Rightarrow Var(5X - 6)=25\times 3=75"